A company just paid a dividend of $2.80 with a growth rate of 2.5% for 10 years. What is the stock price in 6 years if the expected return is 16%?

Thanks in advanced!
Answer #1

Stock price in 6 years (P6) D7=(r-g) Here, =2.8*(1+2.5%)^7 Expected dividend (D7) Required return (r) Growth rate (g) 3.33 16

*Please rate thumbs up